V1 : 3.5V RS: 7500 RL : 20

V1 : 3.5V

RS: 7500

RL : 20
Answer #1

256 Rs 870 V la 4-1 VF + lo Ib RL 50 HGK ca) Assume ideol. opamp la = 16=0 V - Vit. 16 (15) = 16.x5=4V 16+4 20 V. V+ = 4V By>> V. = 17.333 - 11. 666 Vo = 5.6667 5.66 < lo operational amplifies in linear region VO 25k 4 5.66 - 66.64 MA 25k Hostalo whNo 10 3.5002 V V, 23.5002, maximum value to reach saturation (e) when switch closed By KCL Vo - 4 tlo No IRL 20 25k 5 66-4 5-